OpenStudy (anonymous):

Can someone please graph this for the poor girl. She'll give you a medal+fan you according to her.

OpenStudy (anonymous):

Lol, you bump your questions by deleteing them and reposting.

OpenStudy (anonymous):

I think thats abuse as it takes priority over the questions of others who have been waiting for there question to be answered.

OpenStudy (anonymous):

omfg! you just like to bother me! -_- you're causing me a headache! -_- im just trying to help this girl! chill!

OpenStudy (yrelhan4):

Ok, what you need to do is... take any two values of x, and find the corresponding values of y.. then plot them on the cartesian plane and join those points. Job done. :)

OpenStudy (anonymous):

Try using google

OpenStudy (yrelhan4):

Now that, is certainly against the COC.

OpenStudy (anonymous):

https://www.desmos.com/calculator/zd9t7smzjk @Thatyoungbuttercup Go to this link.

OpenStudy (anonymous):

I don't like the fact that everybody is up on my question it seems sneaky & odd 0_0. not to be bogus though its just creepy i usually get like 3 people here .-. oh god >.>

OpenStudy (anonymous):

they want u to fan them. greedy people.

OpenStudy (zzr0ck3r):

@Thatyoungbuttercup sorry I didnt get back to you earlier, ok we need to plot this graph y = x/14, the reason for this is because every 14 seconds we will have another person in the world. so y = new people, x = seconds when x =14 seconds 14 seconds have passed and we have one new baby y = 14/14 = 1 so our first point to plot is (14,1) because y=1 when x = 14 at 28 seconds we have 2 more people in the world than we started with. so y = x/14 becomes y = 28/14 = 2, our next point is (28,2) because y = 2 when x = 28 does this make sense?

OpenStudy (zzr0ck3r):

there are people here because its slow and they are bored.

OpenStudy (anonymous):

Where do babies come from

OpenStudy (anonymous):

ask ur mommy and daddy

OpenStudy (zzr0ck3r):

see ^^^^ bored

OpenStudy (anonymous):

well then ask their mommy and daddy

OpenStudy (anonymous):

lmfaooooooooo!!!!!!!!!! anyways lol yeah i kinda get you so this is what agent0smith gave her though idk if its right.

OpenStudy (zzr0ck3r):

im sure it is

OpenStudy (zzr0ck3r):

gave her?

OpenStudy (nincompoop):

there's the stork theory

OpenStudy (anonymous):

@jack111 if u dont have those either than put the thing in wolfram alpha. Mommy + Daddy = ? See what u get. @zzr0ck3r

OpenStudy (nincompoop):

mommy + daddy = potential divorce

OpenStudy (zzr0ck3r):

ok , notice if you look on that graph, and go to x = 28, the y coordinate is 2. thus the point (28,2), so do you understand how to graph it now?

OpenStudy (zzr0ck3r):

you would just put in a bunch of x values and get the y values and then plot that point.

OpenStudy (nincompoop):

lmao @Jack199

OpenStudy (anonymous):

0_o i just want a graph for the poor girl lol & kinda wait but is that graph kinda right or not though?

OpenStudy (anonymous):

yes that's right @nincompoop the mommy beta function and the daddy zeta function are used to model the potential divorce difference in electrical circuits. exactly my point..

OpenStudy (zzr0ck3r):

its perfect

OpenStudy (anonymous):

thanks! @zzr0ck3r !:D & i get you now though!

OpenStudy (zzr0ck3r):

@Thatyoungbuttercup go to the website wolframalpha.com type graph y = x/14 hit enter

OpenStudy (anonymous):

lol nice one @Jack199

OpenStudy (anonymous):

already did it wait but would it matter what graph i pick? so wait let me get this straight both graphs are good so she can used any one right?

OpenStudy (zzr0ck3r):

wow 15 people in here and the question right below it has no response @genius12 @Jack199 @nincompoop you all should be ashamed. I guess priscilla needs to change her profile pic?

OpenStudy (anonymous):

Wolframalpha says there are roughly 480,000 people alive with the name Jack

OpenStudy (anonymous):

lmfaooooooo!!!!!!! i change my profile pic today .-. i dont think that counts lol & i feel bad for priscilla </3 man where is @mary.rojas when i need her .-.

OpenStudy (anonymous):

& everybody is back on here lol

OpenStudy (anonymous):

k nevermind this is death.

OpenStudy (usukidoll):

eh?

OpenStudy (anonymous):

@zzr0ck3r & everybody else delete your comments!

OpenStudy (zzr0ck3r):

why?

OpenStudy (usukidoll):

wow I only commented once on here.

OpenStudy (anonymous):

for real though delted them!

OpenStudy (usukidoll):

or else what? It's not like I'm doing anything wrong here. Get a mod...they'll find nothing.